the smaller number is 10 less than the larger. the sum of the squares of the numbers is 250. find the numbers
Answer by ewatrrr(24785)


Hi,
Let (x-10) and x represent the smaller and larger number respectively
Question States***
(x-10)^2 + x^2 = 250
Solving for x
x^2 - 20x + 100 + x^2 = 250
2x^2 -20x -150 = 0
2(x^2 - 10x - 75) = 0
factoring
(x-15)(x +5) = 0 Note:SUM of the inner product(-15*x) and the outer product(5x) = -10x
(x-15)= 0
x = 15, the larger number,smaller number is 5
(x +5) = 0
x = -5, the larger number, -15 the smaller number
CHECKING our Answers
|smaller|^2 + |larger|^2 = 250